Bill's gas tank is 1/4 full. After he buys 9 gallons of gas, it is 5/8 full. How many gallons can Bill's tank hold?

Answer by mananth(16949)

let tank hold x gallons
..
it is 1/4 full
so = x/4
added 9 gallons
x/4 +9
= 5/8 of tank
= 5x/8
x/4 +9 = 5x/8
x/4-5x/8 =-9
(2x-5x)/8 = -9
-3x=-72
/-3
x=24 gallons
...
m.ananth@hotmail.ca